OpenStudy (anonymous):

simplify −3x√(20x^2)

OpenStudy (abb0t):

solved.

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

Oh hey, can you try my problem too?

OpenStudy (anonymous):

simplify \[-3x \sqrt{20x ^{2}}\]

OpenStudy (abb0t):

\[\sqrt{5 \times (4\times4) (x\times x})\]

OpenStudy (abb0t):

sorry i meant \(2 \times 2\) not \(4 \times 4\)

OpenStudy (anonymous):

-6 sqrt(5) x sqrt(x^2)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i= −3x√%2820x%5E2%29+ like what Peasant1 said